6 gal of a 28% saline solution was mixed with 3 gal of a 82% solution. Find the concentration of the new mixture.
Answer by Alan3354(69443)

0.28*6 + 0.82*3 = c*9
9c = 1.68 + 2.46 = 4.14
c = 0.46 = 46%